Given that  μX = 3 and  σ2 over X = 16,

a) Let Y = 7X + 43 and find σ2 over Y: ___________ , and μY  : ___________

b) Let Z = 7X + 60 and find  σ2 over z : ________________ , and  μZ : _______________
Answer #1

Solution-A:

mean(Y)=E(Y)=7*E(X)+43=7*3+43=64

μY =64

var(7X+43)=7^2*var(x)=7^2*16=784

ANSWER:

σ2=784

μY =64

Solution-B:

μZ=E(7X+60=7*E(X)+60=7*3+60=81

var(7X+60)=7^2*(var(X)=49*16=784

ANSWER:

σ2=784

μZ=81